[Brahms-l] Shift report 20040422 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Thu Apr 22 2004 - 15:42:47 EDT
------------------------------------------------------------
Date: Thu Apr 22 15:42:46 2004, Shift: 08:00-16:00
Supervisor: BB, with: 
Log book#15 pages 108 - 108

Shift Summary:
------------------------------------------------------------
1. No beam for entire shift

2. Pedestal runs

3. D3 was on 1/2 field for an hour, then tripped.
   The contactor coils in the power supply for D3
   were replaced. Magnet holding steady at 1/2
   field for ~2 hours.

   Problems with DAQ hangups were cured. See comment 
   in e-Log.

   A light leak is apparent in tube #17 in the right
   array. Its high voltage has been reduced to zero.



------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Apr 22 15:43:02 2004

This archive was generated by hypermail 2.1.8 : Thu Apr 22 2004 - 15:43:26 EDT